OpenStudy (anonymous):

Write the equation for the line passing through (5,4) and perpendicular to y=1/8 x+3

OpenStudy (anonymous):

y=-8b+44

OpenStudy (angela210793):

perpendicular----> m=-1/m1=-1/(1/8)=-8 y-y1=m(x-x1) y-4=-8(x-5) y=-8x+44